OpenStudy (anonymous):

what's 1/5 of a mile in feet?

OpenStudy (anonymous):

1056 Ft.

OpenStudy (anonymous):

Your welcome

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

1320

OpenStudy (texaschic101):

1 mile = 5280 1/5(5280) = 5280/5 = 1056 ft

OpenStudy (texaschic101):

oops..1 mile = 5280 ft

OpenStudy (anonymous):

she wants 1/4 a mile

OpenStudy (anonymous):

oh nvm

OpenStudy (anonymous):

lol

OpenStudy (texaschic101):

well..if she wants 1/4th of a mile, then your answer is correct...1320 ft

OpenStudy (anonymous):

yeah but she needs 1/5 a mile. did not see that

OpenStudy (texaschic101):

don't worry...if I had $1 for every mistake I made, I would be rich :)

OpenStudy (amistre64):

the key is having to pay $1 for each mistake ;)

OpenStudy (anonymous):

lol